OpenStudy (anonymous):

Per single gallon of gas, Gina's vehicle can go 16 more miles than Amanda's vehicle. If the combined distance the vehicles can travel on one gallon of gas is 72 miles, what is the distance that Amanda's vehicle travels?

OpenStudy (alekos):

G=A+16 and G+A=72 so (A+16) +A = 72 hence 2A+16=72 2A = 56 A = 28